OpenStudy (volleyballlover55):

Triangle PQR is transformed to triangle P'Q'R'. Triangle PQR has vertices P(8, 0), Q(6, 2), and R(−2, −4). Triangle P'Q'R' has vertices P'(4, 0), Q'(3, 1), and R'(−1, −2). Plot triangles PQR and P'Q'R' on your own coordinate grid. Part A: What is the scale factor of the dilation that transforms triangle PQR to triangle P'Q'R'? Explain your answer. (4 points) Part B: Write the coordinates of triangle P"Q"R" obtained after P'Q'R' is reflected about the y-axis. (4 points) Part C: Are the two triangles PQR and P''Q''R'' congruent? Explain your answer. (2 points)

OpenStudy (volleyballlover55):

@Michele_Laino

OpenStudy (michele_laino):

first step we have to compute the distances PQ and P'Q'

OpenStudy (volleyballlover55):

ok..

OpenStudy (michele_laino):

we have this: \[PQ = \sqrt {{{\left( {8 - 6} \right)}^2} + {{\left( {0 - 2} \right)}^2}} = ...\]

OpenStudy (volleyballlover55):

ok..

OpenStudy (michele_laino):

please complete the computation

OpenStudy (volleyballlover55):

ok ill try :)

OpenStudy (michele_laino):

ok! :)

OpenStudy (volleyballlover55):

PQ= \[\sqrt{8}\]

OpenStudy (michele_laino):

that's right! we get this: \[PQ = \sqrt {{{\left( {8 - 6} \right)}^2} + {{\left( {0 - 2} \right)}^2}} = \sqrt {4 + 4} = \sqrt 8 = 2\sqrt 2 \]

OpenStudy (volleyballlover55):

ok!

OpenStudy (michele_laino):

next, we have: \[P'Q' = \sqrt {{{\left( {4 - 3} \right)}^2} + {{\left( {0 - 1} \right)}^2}} = ...\] again, please complete the computation

OpenStudy (volleyballlover55):

ok!

OpenStudy (volleyballlover55):

P'Q'=2

OpenStudy (michele_laino):

are you sure?

OpenStudy (volleyballlover55):

yes the square root of 2

OpenStudy (michele_laino):

that's right! \[P'Q' = \sqrt {{{\left( {4 - 3} \right)}^2} + {{\left( {0 - 1} \right)}^2}} = \sqrt 2 \]

OpenStudy (michele_laino):

now, the dilation factor, is defined as the ratio between the lengths of corresponding sides, so we can write: \[scale\;factor = \frac{{PQ}}{{P'Q'}} = \frac{{2\sqrt 2 }}{{\sqrt 2 }} = ...\]

OpenStudy (michele_laino):

please simplify

OpenStudy (volleyballlover55):

ok let me try..:)

OpenStudy (volleyballlover55):

2.8 / 1.4?

OpenStudy (michele_laino):

it is simple: